# Exploring Variance Reduction in Importance Sampling for Efficient DNN Training

### v1

@@ 1 Introduction
On the basis of the proposed method, this paper also proposes an effective minibatch size (EMS) for automatic learning-rate adjustment. We derive the EMS $N_{\mathrm{ems}}$ for importance sampling with a minibatch size of $N$ such that uniform sampling with a minibatch size of $N_{\mathrm{ems}}$ achieves the same variance as that obtained with importance sampling. By leveraging the relationship between minibatch size and optimal learning rate [25, 26, 18], we derived an automatic learning-rate adjustment based on the EMS. We also introduce an absolute metric to evaluate the efficiency of importance sampling and designed an importance-score-estimation algorithm: The metric assigns a value of 00 to indicate the theoretically optimal case and $1$ to represent equivalence with uniform sampling. With this metric, the algorithm estimates the importance score of each data point during training by using the moving statistics of per-sample loss gradients. Specifically, the hyperparameter for the moving statistics, which determines the weight assigned to past observations, is designed on the basis of this metric.
@@ 2 Related work
Several studies investigated the use of importance sampling to improve the efficiency of DNN training. Alain et al. [1] proposed a method for distributed training with importance sampling. With this method, the gradient norm of each training sample is computed across multiple computation nodes, aggregated on a master node, and used to carry out training with importance sampling. Although they discussed the variance reduction achieved through importance sampling, they did not explore efficient estimation methods or applications such as learning-rate adjustment, as proposed in this paper. Johnson and Guestrin [11] investigated accelerating SGD training by analyzing the convergence speed of SGD to the optimal solution, leading to the development of importance sampling based on per-sample gradient norms. They proposed a robust regression model to estimate the importance score of each training sample. They also introduced a method for automatically adjusting the learning rate based on the estimated ”expected squared norm of the gradient.” This method differs from the EMS-based learning-rate adjustment we derived. Similarly, Katharopoulos and Fleuret [14] proposed an importance-sampling method for accelerating the convergence speed of SGD to the optimal solution. They introduced a two-step approach for constructing minibatches with importance sampling. Importance scores are first calculated for a large initial minibatch (larger than the final minibatch size), then, a subsampling step is carried out on the basis of the computed importance scores to generate the final minibatch. Katharopoulos and Fleuret [13] proposed using per-sample loss values as importance scores. To speed up the estimation of these loss values, they suggested training a smaller auxiliary model for loss estimation alongside the primary model. Chang et al. [4] proposed a method for maintaining a history of model predictions during training and estimating per-sample importance scores using either the average or variance of these predictions.
@@ 3.1 Notations
Let the training dataset, consisting of $M$ pairs of input $x$ and ground truth label $y$, be represented as $\mathcal{D}_{\mathrm{train}}=\left\{\left(x^{(i)},y^{(i)}\right)\right\}_{i=1}^{M}$. Each pair $\left(x^{(i)},y^{(i)}\right)$ is assumed to be i.i.d. in accordance with the distribution $p(x,y)$. Let $f_{\theta}$ denote a DNN model to be trained, where $\theta$ represents the model parameters. The training loss function for the $i$-th data point $\left(x^{(i)},y^{(i)}\right)$ is defined as $\mathcal{L}(\theta;i):=\ell\left(f_{\theta}\left(x^{(i)}\right),y^{(i)}\right)$, where $\ell$ can be a function such as the cross-entropy function. The gradient of the loss function with respect to $\theta$ is denoted as $\nabla_{\theta}\mathcal{L}(\theta;i)$. Since $\theta$ represents a vector, $\nabla_{\theta}\mathcal{L}(\theta;i)$ is also a vector. Specifically, the gradient with respect to the $k$-th parameter $\theta_{k}$ is represented as $\nabla_{\theta_{k}}\mathcal{L}(\theta;i)$, which is a scalar. Additionally, for a natural number $a$, let $\llbracket a\rrbracket:=\left\{1,\ldots,a\right\}$. The expectation and variance (or covariance matrix) of a random variable $z$ following the distribution $p(z)$ are denoted as $\mathbb{E}_{z\sim p(z)}\left[\cdot\right]$ and $\mathbb{V}_{z\sim p(z)}\left[\cdot\right]$, respectively. Note that these become a vector and a matrix, respectively, when $z$ is a vector. Let >$\mathbb{R}^{M}_{>0}$ denote the $M$-dimensional space in which all elements are strictly positive.
@@ 3.2 DNN training based on uniform sampling
In DNN training based on uniform sampling, the following expectation of the loss gradient is computed:
$\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]:=\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\nabla_{\theta}\mathcal{L}(\theta;i)\right],$ (1)
where $p_{\mathrm{unif}}(i)$ denotes a uniform distribution defined over $\llbracket M\rrbracket$, the index set of $\mathcal{D}_{\mathrm{train}}$, indicating $p_{\mathrm{unif}}(i)=1/M\ \left(\forall i\in\llbracket M\rrbracket\right)$. Since $\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\nabla_{\theta}\mathcal{L}(\theta;i)\right]\approx\mathbb{E}_{(x,y)\sim p(x,y)}\left[\nabla_{\theta}\ell\left(f_{\theta}\left(x\right),y\right)\right]$, the expectation computed using uniform sampling from $\mathcal{D}_{\mathrm{train}}$ approximately corresponds to the expectation under $p(x,y)$. In standard SGD training, $\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ is estimated using a minibatch sampled in accordance with $p_{\mathrm{unif}}(i)$. [^2] In SGD, sampling is often carried out without replacement (no resampling) within each epoch. However, for ease of comparison with importance sampling, we assume uniform sampling with replacement. For example, Chang et al. [4] refers to the former as SGD-Scan and the latter as SGD-Uni. Both methods were compared in our experiments.
@@ 3.3 Training with importance sampling
Let $W\in\mathbb{R}^{M}_{>0}$ denote the importance score, where the $i$-th element $w_{i}$ represents the importance of $\left(x^{(i)},y^{(i)}\right)\in\mathcal{D}_{\mathrm{train}}$. In DNN training using importance sampling with $W$, the following expectation is computed:
$\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ $:=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\nabla_{\theta}\mathcal{L}(\theta;i)\right],$ (2)
$r(i;W)$ $:=\frac{p_{\mathrm{unif}}(i)}{p_{\mathrm{is}}(i;W)},$
where $p_{\mathrm{is}}(i;W)$ is a distribution defined over $\llbracket M\rrbracket$ and given by $p_{\mathrm{is}}(i;W)\propto w_{i}$, which indicates $p_{\mathrm{is}}(i;W)=w_{i}/\sum_{i^{\prime}\in\llbracket M\rrbracket}w_{i^{\prime}}$. The term $r(i;W)$ is referred to as the importance-sampling weight, or simply the weight, of data point $i$.
@@ 3.4 Expectation equivalence
The following relationship holds for any $W\in\mathbb{R}^{M}_{>0}$:
$\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]=\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right],$ (3)
which indicates that the expectation of the loss gradient remains identical under both uniform sampling and importance sampling, regardless of $W$ (see subsection A.1 for a detailed derivation of Eq. 3).
@@ 3.5 Comparison of variance
Although the expectations of the loss gradient for uniform sampling and importance sampling are identical regardless of $W$, their variances may vary depending on $W$.
@@
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{\theta_{k}\in\theta}\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right],$
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{\theta_{k}\in\theta}\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right].$
This corresponds to summing the variances of the gradients for each $\theta_{k}$, while ignoring the covariances between parameters. It has been shown [1] that $\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ is minimized when $W$ is set to the following optimal value $W^{\ast}$:
@@ 4 Variance estimation of loss gradient
Proposition 4.1 (Variance estimation for importance sampling).
@@
The key point of Proposition 4.1 is that the expectations in the variance-estimation formulas are all taken with respect to $p_{\mathrm{is}}(i;W)$. This enables the traces of the covariance matrices to be computed using only samples drawn from $p_{\mathrm{is}}(i;W)$. In other words, during training with minibatches under importance sampling, all relevant variances can be estimated solely from the data points within the minibatch.
4.1 Pseudo code
In Appendix B, we present the pseudocode for variance estimation based on Proposition 4.1. As shown in the pseudocode, the traces of the covariance matrices described in Proposition 4.1 can be efficiently estimated with only a few lines of code, given $r(i;W)$ and the loss gradients for each data point in a minibatch generated in accordance with $p_{\mathrm{is}}(i;W)$.
4.2 On computing loss gradient for each data point
The formula in Proposition 4.1 requires the computation of the loss gradient $\nabla_{\theta}\mathcal{L}\left(\theta;i\right)$ for each data point. However, calculating the loss gradient with respect to $\theta$ introduces significant computational overhead. To address this issue, methods have been proposed to use the gradient with respect to the output of the final layer of the DNN (referred to as the logits) instead of the gradient with respect to $\theta$ [11, 14]. Compared with computing the gradient for $\theta$, the gradient with respect to the logits can be obtained at a significantly lower computational cost. The L2 norm of the gradient with respect to the logits is then used as an approximation, up to a constant factor, of the L2 norm of the gradient with respect to $\theta$. Experimental results indicated that these values exhibit a high correlation [11]. In our experiments, we also adopted the gradient with respect to the logits to enable efficient computation.
@@ 5 Effective-minibatch-size estimation and learning-rate adjustment
We derived an EMS that achieves the same variance reduction under uniform sampling as that achieved by importance sampling with a minibatch of size $N$ and $W$. We then propose its application to automatic learning-rate adjustment.
@@ 5.1 Effective minibatch size
In contrast to ESS [15, 19], denoted as $N_{\mathrm{ess}}$, which is implicitly restricted to $N_{\mathrm{ess}}\leq N$ [21], the proposed EMS can satisfy $N_{\mathrm{ems}}>N$ depending on $W$. Therefore, Proposition 5.2 suggests that, with an appropriate $W$, training under importance sampling with a minibatch of size $N$ is equivalent to training under uniform sampling with a minibatch of size $N_{\mathrm{ems}}$, where $N_{\mathrm{ems}}>N$. This suggests that importance sampling with an appropriate $W$ effectively increases the number of training epochs for a fixed number of training iterations, as training with a larger minibatch size corresponds to an increase in the number of effective epochs.

### v2

@@ Abstract
Note. This is the author’s accepted manuscript. The final version is published in SIAM Journal on Mathematics of Data Science and is available via https://doi.org/10.1137/25M1726339.
@@ 1 Introduction
On the basis of the proposed method, this paper also proposes an effective minibatch size (EMS) for automatic learning-rate adjustment. We derive the EMS $N_{\mathrm{ems}}$ for importance sampling with a minibatch size of $N$ such that uniform sampling with a minibatch size of $N_{\mathrm{ems}}$ achieves the same variance as that obtained with importance sampling. By leveraging the relationship between minibatch size and optimal learning rate [30, 31, 21, 26, 22], we derive an automatic learning-rate adjustment based on the EMS. We also introduce an absolute metric to evaluate the efficiency of importance sampling and design an importance-score-estimation algorithm: The metric assigns a value of 0 to indicate the theoretically optimal case and $1$ to represent equivalence with uniform sampling. With this metric, the algorithm estimates the importance score of each data point during training by using the moving statistics of per-sample loss gradients. Specifically, the hyperparameter for the moving statistics, which determines the weight assigned to past observations, is designed on the basis of the proposed efficiency metric.
@@ 2 Related work
Several studies investigated the use of importance sampling to improve the efficiency of DNN training. Alain et al. [1] proposed a method for distributed training with importance sampling. With this method, the gradient norm of each training sample is computed across multiple computation nodes, aggregated on a master node, and used to carry out training with importance sampling. Although they discussed the variance reduction achieved through importance sampling, they did not explore efficient estimation methods or applications such as learning-rate adjustment, as proposed in this paper. Johnson and Guestrin [13] investigated accelerating SGD training by analyzing the convergence speed of SGD to the optimal solution, leading to the development of importance sampling based on per-sample gradient norms. They proposed a robust regression model to estimate the importance score of each training sample. They also introduced a method for automatically adjusting the learning rate based on the estimated “expected squared norm of the gradient.” This method differs from the EMS-based learning-rate adjustment we derived. Similarly, Katharopoulos and Fleuret [16] proposed an importance-sampling method for accelerating the convergence speed of SGD to the optimal solution. They introduced a two-step approach for constructing minibatches with importance sampling. Importance scores are first calculated for a large initial minibatch (larger than the final minibatch size), then, a subsampling step is carried out on the basis of the computed importance scores to generate the final minibatch. Katharopoulos and Fleuret [15] proposed using per-sample loss values as importance scores. To speed up the estimation of these loss values, they suggested training a smaller auxiliary model for loss estimation alongside the primary model. Chang et al. [4] proposed a method for maintaining a history of model predictions during training and estimating per-sample importance scores using either the average or variance of these predictions.
@@
Complementary SGD improvement strategies
Besides importance-sampling-based variance reduction, SGD has been improved through various approaches, ranging from optimizer design (e.g., Adam [17]) to alternative variance-reduction techniques. The latter include control-variate-based methods such as SVRG [11], which reduce gradient variance by periodically computing full-batch gradients or maintaining historical gradient tables. While effective in convex or small-scale settings, these methods often face practical challenges in deep learning due to the breakdown of underlying assumptions [6]. Our approach is orthogonal to both lines of research. As demonstrated in our experiment in subsection 9.5, it can be combined with the adaptive optimizer Adam by appropriately adjusting the learning rate, as discussed in subsection 5.2.
@@ 3.1 Notations
Let the training dataset, consisting of $M$ pairs of input $x$ and ground truth label $y$, be represented as $\mathcal{D}_{\mathrm{train}}=\left\{\left(x^{(i)},y^{(i)}\right)\right\}_{i=1}^{M}$. Each pair $\left(x^{(i)},y^{(i)}\right)$ is assumed to be independent and identically distributed (i.i.d.) in accordance with the distribution $p(x,y)$. Let $f_{\theta}$ denote a DNN model to be trained, where $\theta=\left(\theta_{k}\right)_{k=1}^{K}$ represents the model parameters. The training loss function for the $i$-th data point $\left(x^{(i)},y^{(i)}\right)$ is defined as $\mathcal{L}(\theta;i):=\ell\left(f_{\theta}\left(x^{(i)}\right),y^{(i)}\right)$, where $\ell$ is assumed to be a continuously differentiable function, such as the cross-entropy loss. The gradient of the loss function for the $i$-th data point with respect to $\theta$ is denoted as $\nabla_{\theta}\mathcal{L}(\theta;i)$. Since $\theta$ represents a vector, $\nabla_{\theta}\mathcal{L}(\theta;i)$ is also a vector. Specifically, the gradient with respect to the $k$-th parameter $\theta_{k}$ is represented as $\nabla_{\theta_{k}}\mathcal{L}(\theta;i):=\frac{\partial\mathcal{L}(\theta;i)}{\partial\theta_{k}}$, which is a scalar. Additionally, for a natural number $a$, let $\llbracket a\rrbracket:=\left\{1,\ldots,a\right\}$. The expectation and variance (or covariance matrix) of a random variable $z$ following the distribution $p(z)$ are denoted as $\mathbb{E}_{z\sim p(z)}\left[\cdot\right]$ and $\mathbb{V}_{z\sim p(z)}\left[\cdot\right]$, respectively. Note that these become a vector and a matrix, respectively, when $z$ is a vector. Let $\mathbb{R}^{M}_{>0}$ denote the $M$-dimensional space in which all elements are strictly positive. Let $\|\cdot\|$ denote the L2 norm of a vector.
@@ 3.2 DNN training based on uniform sampling
At each training step, we would like to know the true gradient $G(\theta):=\mathbb{E}_{(x,y)\sim p(x,y)}\left[\nabla_{\theta}\ell\left(f_{\theta}\left(x\right),y\right)\right]$, which we don’t have access to. The goal of any gradient estimator is to approximate $G(\theta)$ as well as possible with (a subset of) the training data. In DNN training based on uniform sampling, the gradient of the loss is estimated by $\nabla_{\theta}\mathcal{L}(\theta;i)$ with $i\sim p_{\mathrm{unif}}(i)$, where $p_{\mathrm{unif}}(i)$ denotes a uniform distribution defined over $\llbracket M\rrbracket$, the index set of $\mathcal{D}_{\mathrm{train}}$, i.e., $p_{\mathrm{unif}}(i)=1/M\ \left(\forall i\in\llbracket M\rrbracket\right)$. We consider the following expectation:
$\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]:=\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\nabla_{\theta}\mathcal{L}(\theta;i)\right],$
where $\mathcal{L}(\theta)$ in the left hand side is used as a shorthand for the per-sample loss $\mathcal{L}(\theta;i)$, and $\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ represents the expectation of the per-sample gradient under uniform sampling. From the assumption of each $(x^{(i)},y^{(i)})\in\mathcal{D}_{\mathrm{train}}$ being i.i.d. with respect to $p(x,y)$, we have $G(\theta)\approx\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$. In standard SGD training, we approximate $\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ by computing the average gradient over a minibatch sampled according to $p_{\mathrm{unif}}(i)$. [^2] In SGD, sampling is often carried out without replacement (no resampling) within each epoch. However, for ease of comparison with importance sampling, we assume uniform sampling *with* replacement. For example, Chang et al. [4] refer to the former as SGD-Scan and the latter as SGD-Uni. Both methods were compared in our experiments. This minibatch gradient is an unbiased estimate of $G(\theta)$.
@@ 3.3 Training with importance sampling
Let $W\in\mathbb{R}^{M}_{>0}$ denote the vector of importance weights, where the $i$-th element $w_{i}$ indicates the importance of $(x^{(i)},y^{(i)})\in\mathcal{D}_{\mathrm{train}}$. In DNN training with importance sampling based on $W$, the loss gradient is estimated as $r(i;W)\,\nabla_{\theta}\mathcal{L}(\theta;i)$, where the index $i$ is drawn from the distribution $p_{\mathrm{is}}(i;W)$ defined over $\llbracket M\rrbracket$ by
$p_{\mathrm{is}}(i;W):=\frac{w_{i}}{\sum_{i^{\prime}\in\llbracket M\rrbracket}w_{i^{\prime}}}.$ (1)
This indicates $p_{\mathrm{is}}(i;W)\propto w_{i}$. The corresponding importance-sampling coefficient $r(i;W)$ is given by
$r(i;W):=\frac{p_{\mathrm{unif}}(i)}{p_{\mathrm{is}}(i;W)}.$ (2)
We then define the weighted expectation as
$\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ $:=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\nabla_{\theta}\mathcal{L}(\theta;i)\right].$ (3)
As shown in the next section, $\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]=\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ holds. Therefore, estimating $\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ using a minibatch sampled from $p_{\mathrm{is}}(i;W)$ provides an unbiased estimate of $G(\theta)$.
@@ 3.4 Expectation Equivalence
The following relationship holds between the expected loss gradient under uniform sampling and that under importance sampling.
Proposition 3.1.
For any $W\in\mathbb{R}^{M}_{>0}$, it holds that
$\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]=\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right].$ (4)
Proof.
See subsection A.1. ∎
This proposition indicates that the expected loss gradient remains the same under both uniform sampling and importance sampling, regardless of $W$.
@@ 3.5 Comparison of variance
Although the expectations of the loss gradient for uniform sampling and importance sampling are identical regardless of $W$, as discussed above, their variances may vary depending on $W$.
@@
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{k=1}^{K}\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right],$
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{k=1}^{K}\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right].$
This corresponds to summing the variances of the gradients for each $\theta_{k}$, while ignoring the covariances between parameters. The following result is known regarding the optimal importance sampling weight [1].
Proposition 3.2 (Optimal importance sampling weight [1]).
The importance sampling weight $W^{\ast}$ that minimizes the trace of the gradient variance $\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ is given by
@@
Considering that $p_{\mathrm{unif}}(i)=1/M$ and is independent of $i$, it follows from Eq. 5 that minimizing the trace of the covariance matrix of the loss gradient can be achieved by performing importance sampling with weights proportional to the L2 norm of the loss gradient $\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|$ for each data point. If the L2 norm of the loss gradient serves as a measure of the degree to which a data point $i$ affects the training, prioritizing the sampling of data points with larger effects reduces the variance of the loss gradient.
@@ 4.1 Trace formulas for gradient variance
Proposition 4.1 (Trace of gradient variance under uniform and importance sampling).
@@
4.2 Online variance estimation during minibatch-based training
The key point of Proposition 4.1 is that all expectations in the formulas are taken with respect to $p_{\mathrm{is}}(i;W)$. This allows the traces of the covariance matrices to be estimated using only samples drawn from $p_{\mathrm{is}}(i;W)$. Therefore, during training with minibatches generated via importance sampling, all relevant trace values can be estimated solely from the loss gradients of the data points within the minibatch. For reference, Appendix B presents the pseudocode for variance estimation based on Proposition 4.1, given $r(i;W)$ and the loss gradients of each data point in a minibatch generated according to $p_{\mathrm{is}}(i;W)$.
@@ 5 Effective-minibatch-size estimation and learning-rate adjustment
We derived an EMS that achieves the same variance reduction under uniform sampling as that achieved by importance sampling with importance weights $W$ and a minibatch of size $N$. We then propose its application to automatic learning-rate adjustment.
@@ 5.1 Effective minibatch size
In contrast to ESS [18, 23], denoted as $N_{\mathrm{ess}}$, which is implicitly restricted to $N_{\mathrm{ess}}\leq N$ [25], the proposed EMS can satisfy $N_{\mathrm{ems}}>N$ depending on $W$. Therefore, Proposition 5.2 suggests that, with an appropriate $W$, training under importance sampling with a minibatch of size $N$ is first-order equivalent—in the sense that it matches both the expected stochastic gradient and the trace of its covariance matrix—to training under uniform sampling with a minibatch of size $N_{\mathrm{ems}}$, where $N_{\mathrm{ems}}>N$. This suggests that importance sampling with an appropriate $W$ effectively increases the number of training epochs for a fixed number of training iterations, as training with a larger minibatch size corresponds to an increase in the number of effective epochs.